Consider the following model of the economy: C=C0 +c1Yd
T=t0 +t1Y
Yd = Y - T
G and I are both constant (exogenous).
1
Expert's answer
2020-04-20T18:46:30-0400

Y=C+I+G+Xn

Y=C0+MPC(Y-T0)+I+G+Xn

Y=C-c1(Y-t0-tY)+MPC(Y-t0-t1Y)+I+G+Xn